OpenStudy (anonymous):

Compute the derivative f'(a) algebraically. f(x)=x-2x^3

OpenStudy (tkhunny):

Set up your ratio and have at it. Let's see your best work.

OpenStudy (anonymous):

Remember how to find the slope of a line? \[\frac{y_2 - y_1}{x_2 - x_1}\] The derivative formula is just a fancier version of that. Since \(y = f(x)\) we can say: \[f'(x) = \lim_{x \rightarrow n} \frac{f(x) - f(n)}{x - n}\]